a. A set of n = 15 pairs of X and Y scores has SSx = 10, SSy = 40, and SP = 30. What is the slope for the regression equation for predicting Y from X?

b.  For the regression equation Y = 2X – 3, if the mean for X is x̅ = 5, what is the mean for Y?

Show work
Answer #1

Ansa:

slope=sp/sxx=30/10=3

#slope=3

Ansb:

regression equation  Y = b0+b1*x here b0=-3 and b1=2   x̅ = 5   

b0= -b1* x̅

=b0+b1* x̅

=-3+2*5

=7

#mean of y=7